# Circumstellar Interaction in the Ultraviolet Spectra of SN 2023ixf 14–66 Days After Explosion

### v1

@@ Abstract
SN 2023ixf was discovered in M101 within a day of explosion and rapidly classified as a Type II supernova with flash features. Here we present ultraviolet (UV) spectra obtained with the Hubble Space Telescope 14, 19, 24, and 66 days after explosion. Interaction between the supernova ejecta and circumstellar material (CSM) is seen in the UV throughout our observations in the flux of the first three epochs and asymmetric Mg II emission at day 66. We compare our observations to CMFGEN supernova models which include CSM interaction ($\dot{M}<10^{-3}$ $\mathrm{M_{\odot}~{}yr^{-1}}$) and find that the power from CSM interaction is decreasing with time, from $L_{\rm sh}\approx 5\times 10^{42}$ $\mathrm{erg~{}s^{-1}}$ at day 14 to $L_{\rm sh}\approx 1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$ at day 66. We examine the contribution of individual atomic species to the spectra at day 14 and 19, showing that the majority of the features are dominated by iron, nickel, magnesium, and chromium absorption in the ejecta. The UV spectral energy distribution of SN 2023ixf sits between that of supernovae which show no definitive signs of CSM interaction and those with persistent signatures assuming the same progenitor radius and metallicity. Finally, we show that the evolution and asymmetric shape of the Mg II $\lambda\lambda$2796, 2802 emission are not unique to SN 2023ixf. These observations add to the early measurements of dense, confined CSM interaction, tracing the mass-loss history of SN 2023ixf to $\sim 33$ yr prior to explosion and the density profile to a radius of $\sim 5.7\times 10^{15}$ cm. They show the relatively short evolution from quiescent red supergiant wind to high mass loss.
Type II supernovae; Red supergiant stars; Stellar mass loss; Ultraviolet transient source; Ultraviolet spectroscopy
@@ 1 Introduction
Hydrogen-rich supernovae (Type II) are thought to come from red supergiants (RSGs) with zero-age main sequence (ZAMS) masses between $\sim 8$ and $\sim 25$ $\mathrm{M_{\odot}}$. While it is accepted that all RSGs lose mass, there are no quantitative mass-loss rates derived from stellar evolution theory. Adding to this, there is no consensus about the empirical mass-loss rates for RSGs; Massey et al. (2023) infer mass-loss rates higher than the canonical rates of de Jager et al. (1988) from the RSG luminosity function, while other studies, such as (Beasor et al., 2020), measure lower mass-loss rates from the infrared (IR) excess of RSGs. Different mass-loss rates affect the final star that explodes, impacting the final hydrogen envelope mass as well as the density structure of the circumstellar material (CSM) into which the supernova explodes.
@@
Observational evidence supporting these predictions is unclear. With sample analyses finding contradictory results about whether supernovae with flash features are systematically bluer and more luminous than non-interacitng supernovae (Khazov et al., 2016; Bruch et al., 2023). Additional contradictions are found in individual objects (e.g. SN 2018zd, Hiramatsu et al. 2021; SN 2020pni, Terreran et al. 2022; various, Jacobson-Galán et al. 2024; PTF11iqb, Smith et al. 2015; SN 2013fs, Bullivant et al. 2018). Spectroscopically, while an absorption feature is often observed blueward of $\mathrm{H\alpha}$ (Gutiérrez et al., 2017) and has been interpreted as high-velocity hydrogen due to CSM interaction, the relationship between the sample of supernovae with this feature and those with flash features is unclear. UV observations of Type II supernovae are especially sparse, with early-time data limited by the availability of sensitive rapid-response UV telescopes, and with late-time observations limited by the fading of the UV when CSM interaction is not present.
@@
The study of Type II supernovae with dense, confined CSM was recently enriched by the explosion of SN 2023ixf in the nearby galaxy M101. The proximity, early discovery and classification (Itagaki, 2023; Perley & Gal-Yam, 2023), and identification of narrow emission lines have led to this being the best-observed supernova with flash features to date. Of particular note is the periodic pre-explosion light curve of the RSG progenitor in the optical and IR (Jencson et al., 2023; Kilpatrick et al., 2023; Niu et al., 2023; Qin et al., 2023; Soraisam et al., 2023; Xiang et al., 2024), the incredibly deep multiband pre-discovery data which characterize the light-curve rise starting at an absolute magnitude of $V\approx-11$ together with high-cadence photometry through the end of the plateau phase (Jacobson-Galán et al., 2023; Hiramatsu et al., 2023; Hosseinzadeh et al., 2023; Sgro et al., 2023; Teja et al., 2023; Yamanaka et al., 2023; Zimmerman et al., 2024; Bersten et al., 2024; Martinez et al., 2024), intranight optical spectroscopy for more than a week after discovery (Bostroem et al., 2023a; Jacobson-Galán et al., 2023; Teja et al., 2023; Hiramatsu et al., 2023; Yamanaka et al., 2023; Zhang et al., 2023), early-time optical spectropolarimetry (Vasylyev et al., 2023a; Singh et al., 2024), high-resolution spectroscopy of the first 17 days of evolution (Smith et al., 2023; Hiramatsu et al., 2023; Zimmerman et al., 2024), X-ray and radio detections (Berger et al., 2023; Grefenstette et al., 2023; Matthews et al., 2023; Chandra et al., 2024), and the first early-time UV spectra of a supernova with flash features (Zimmerman et al., 2024). While a coherent picture of these observations is still being explored, they point toward a confined, dense CSM with a radius $R\lesssim 10^{15}$ cm and a mass-loss rate $\dot{M}\approx 10^{-2}$$10^{-4}$ $\mathrm{M_{\odot}~{}yr^{-1}}$.
@@ 2 Observations and Data Reduction
HST observations were obtained on days 14, 19, 24, and 66 with the CCD detector of the Space Telescope Imaging Spectrograph (STIS) using the 52″$\times$ 0.2″slit at the E1 position to mitigate flux loss to charge transfer efficiency loss and the G230LB, G430L, and G750L gratings (DD GO-17313; Bostroem et al., 2023b). Unfortunately, our planned visit on day 50 failed to acquire guide stars. However, the observation was successfully repeated on day 66. We obtained at least 4 exposures with each grating to automatically remove cosmic rays. A list of observations is given in Table 1.
@@
The G750L grating suffers fringing at the reddest wavelengths which can be corrected using a contemporaneous fringe flat. We obtained a fringe flat with each visit using the $0.3^{\prime\prime}\times 0.09^{\prime\prime}$ aperture, as recommended for the E1 aperture position. The fringe flat was applied to the G750L observations using the module stistools.defringe and 1D spectra were extracted using the stistools.x1d routine.
We utilized data-quality flags 16 (high dark rate) and 512 (bad reference pixel) to eliminate bad pixels. In particular, we note that narrow emission features that appear in all spectra are flagged as high dark rate pixels and therefore removed from the analysis. We confirm that these are isolated pixels with high count rates in the dark files by examining the extraction location in the 2D dark images themselves. Nevertheless, it is possible that there is a real unresolved emission line at the same pixel as at least one of these features. However, this would be contaminated by the high dark rate and therefore unable to provide any constraint on the CSM. Thus, given the characteristics described above, we treat these features as artifacts and do not consider them further.
@@
We add to this dataset the HST observations taken between days 3 and 11 as part of GO-17205 (Zimmerman et al., 2022) which have been presented by Zimmerman et al. (2024). 1D-extracted and flux-calibrated files were downloaded from the Mikulski Archive for Space Telescopes. As noted by Zimmerman et al. (2024), some of these exposures suffered from pointing, acquisition, and/or saturation issues which led to unreliable flux due to the supernova not being aligned in the slit and the nonlinearity of the CCD response near the saturation point. To mitigate these issues, we scale all observations with each grating in a given visit to the spectrum with the most flux (assuming this is the best-centered observation) using a first-degree polynomial fit to the ratio of the spectra as the scale factor. A constant multiple of a few percent was then applied to align the G430L and G750L observations to the G230LB observation, and then the fluxes were combined by taking the median value at each wavelength, excluding pixels with data-quality flags of 16 and 512. We note that the majority of pixels are saturated in the range $\sim 3200$–5000 Å in visit 1, $\sim 2100$–7600 Å in visits 2 and 3, and $\sim 2900$–7900 Å in visit 4, making these wavelength ranges the most unreliable in terms of flux calibration. However, in this paper we are primarily analyzing the later epochs ($14-66$ d) which are not saturated and only use the early epochs ($<10$ d) to identify the presence and timing of relevant spectral features. Thus, we do not attempt to further correct for this effect.
@@ 3 Spectral Evolution
The near-UV (NUV) through near-infrared (NIR) evolution of SN 2023ixf from day 3 to 66 is shown in Figure 1. No Type II supernovae (excluding Type IIn supernovae) since SN 1987A have been observed this early in the UV and with this cadence, revealing unprecedented details of the evolution. As has been noted by other authors, narrow emission lines are present in the early-time spectra, including notably a narrow P Cygni line from N IV $\lambda 1718$ which is present in the first spectrum and fades by day 8, and C III $\lambda 2297$ which is present from day 4 to day 8 (Yamanaka et al., 2023; Jacobson-Galán et al., 2023; Bostroem et al., 2023a; Teja et al., 2023; Hiramatsu et al., 2023; Zimmerman et al., 2024). These early narrow lines indicate dense CSM surrounding the supernova which has not been shocked. By day 14, these lines disappear and the optical spectrum is that of a young Type II supernova. Although slow to develop as a result of early interaction, a typical Type II supernova optical spectrum emerges with prominent Doppler broadened P Cygni profiles in H I, He I, and other metals.
@@
Figure 1: The evolution of the NUV to NIR spectra of SN 2023ixf from 3.5 to 66 days after explosion. Prominent features (N IV: 1718, 7109, 7122 Å; C III: 2297 Å; Mg II: 2796, 2802 Å; He II: 3203, 4685.5, 4860 Å; C IV: 5801, 5812 Å; $\mathrm{H\alpha}$: 6563 Å; He I: 6678.1 Å; Ca II: 8498, 8542, 8662 Å) are marked with vertical lines at their rest wavelength and labeled at the top of the figure. Narrow ISM absorption lines have been removed to highlight the supernova spectra.
@@
A more detailed view of the evolution from day 14 through day 66 is shown in Figure 2. Although still dominated by iron line blanketing (the broad absorption of flux due to a forest of iron lines which blend together), the UV continuum flux is clearly present in the day 14 spectrum. This is unusual for classical Type II supernovae in which the UV flux fades rapidly (Baron et al., 2000; Brown et al., 2007; Vasylyev et al., 2023b; Bostroem et al., 2023c).
@@
Figure 2: A comparison of the spectra of SN 2023ixf from 14.25 to 66.25 days post-explosion (solid lines) with the best-matching CSM interaction model (semi-transparent). Over time, the UV emission declines and P Cygni features develop in the optical, until day 66, when only Mg II $\lambda\lambda$2796, 2802 is visible in the UV. This corresponds to a sequence of models with decreasing interaction power. Even though the model spectra do not match the SED of the observations, especially in the first two spectra where they appear to be cooler (resulting in lower-ionization species and more UV line blanketing), the match is impressive given the 1D nature and the ad-hoc implementation of the mass, density, and clumping of the cool dense shell. Prominent spectral features have been marked with dotted lines at 7000 $\mathrm{km~{}s^{-1}}$, the approximate location of the absorption throughout the spectroscopic sequence. The “emission” feature at 1925 Å is marked at rest with a dashed line. Narrow ISM absorption lines have been removed to better highlight the supernova spectra. Observed and model spectra at each epoch have been placed at the same distance with the same offset applied.
@@
A comparison of the Mg II and $\mathrm{H\alpha}$ profiles provides insight into the origin of the Mg II emission (see Figure 3). When CSM is present, a cool dense shell forms at the interface between the ejecta and CSM. While the $\mathrm{H\alpha}$ emission originates from the ejecta and the cool dense shell, the Mg II is thought to originate only from the shell and should thus be at a similar velocity as the outer edges of the ejecta. We compare Mg II and $\mathrm{H\alpha}$ in Figure 3, finding that the profiles extend to similar velocities on both the red and blue sides, confirming that the Mg II is emitted from the cool dense shell at the outer edge of the ejecta. However, $\mathrm{H\alpha}$ sits in the optical where there is continuum and forms throughout the ejecta and cool dense shell, leading to a P Cygni profile, while Mg II, without UV continuum and forming in the outer ejecta, is seen solely in emission. There is a shallow, narrow dip at the blue edge of the $\mathrm{H\alpha}$ profile which could be absorption from the cool dense shell at $\sim 9200$ $\mathrm{km~{}s^{-1}}$, although it is not clear if this feature is real given the signal-to-noise ratio (S/N) of the spectrum.
@@
Figure 3: Left: the $\mathrm{H\alpha}$ and Mg II line profiles at day 66 in velocity space of SN 2023ixf compared to the CSM interaction models of Dessart & Hillier (2022) with $L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$, scaled to the distance of SN 2023ixf. The shape of both lines are well captured by the models, given that the cool dense shell is at a higher velocity in the models than in SN 2023ixf. Right: the Mg II profile (black) fit with a pure Gaussian (pink) modeling spherical ejecta, a symmetric shell (blue dotted) modeling emission from the cool dense shell in the absence of optical-depth effects, and an asymmetric shell (green dashed) modeling the cool dense shell with optical depth.
@@
To characterize the Mg II feature further, we fit three different profiles in Figure 3, representing different physical configurations. A spherical distribution of material with a Gaussian emitting profile will produce a Gaussian profile. Although the profile is clearly not Gaussian, it could be coming from a spherical distribution of material which is being attenuated. To explore the non-attenuated spherical distribution, we fit a Gaussian to the blue side of the Mg II profile from $-10,000$ to $-7000$ $\mathrm{km~{}s^{-1}}$. This fit is shown as a pink line in Figure 3; it has a mean of $\mu=0$ $\mathrm{km~{}s^{-1}}$ and a full width at half-maximum intensity (FWHM) of 7780 $\mathrm{km~{}s^{-1}}$. The observed blue side of Mg II is well described by a Gaussian profile. An alternate emitting region could be a sphere with a Gaussian profile which has had the center removed to form a thick shell. This configuration will have a flat-topped profile with Gaussian wings (Jerkstrand, 2017) if the cool dense shell and ejecta are optically thin. Given our slanted top profile, we use the maximum of our observed flux as the flat-topped maximum of the model profile. This defines the inner radius velocity of the shell as $v_{\rm in}=-6800$ $\mathrm{km~{}s^{-1}}$, and the profile is shown as a dotted-blue line in the right panel of Figure 3.
@@
To model the slated top of the profile, we model the line as if it were generated by an asymmetric shell of material (i.e., a shell in which there is more flux from one side than the other reaching the observer). Following Kwok et al. (2023, 2024), we model this asymmetric shell as a Gaussian profile with an off-center hole cut out which results in a slanted top instead of a flat-topped profile. We parameterize this model by the FWHM and mean of the Gaussian, the velocity of the center of the hole ($v_{\rm c}$), and the velocity of the radius of the hole ($v_{\rm in}$). We find FWHM = 7580 $\mathrm{km~{}s^{-1}}$, $\mu=-2995$ $\mathrm{km~{}s^{-1}}$, $v_{\rm in}=5000$ $\mathrm{km~{}s^{-1}}$, and $v_{\rm c}=1750$ $\mathrm{km~{}s^{-1}}$, where $v_{\rm c}$ and $v_{\rm in}$ are measured relative to $\mu$. This model fits the data exceptionally well and is shown as a green dashed line in the right panel of Figure 3. While the CSM is likely asymmetric, we caution that this parameterization is more likely to indicate different optical depths along the line of sight through the ejecta and cool dense shell, and should not be taken as a literal description of the ejecta. It does, however, demonstrate that this profile can be created by a shell of material in which a higher flux of photons from the blue side of the ejecta is reaching us, than photons from the red side. Additionally, it defines the inner edge of the blue side of the shell to be at velocity $-6245$ $\mathrm{km~{}s^{-1}}$(where the profile peaks), close to our earlier estimate of $-6800$ $\mathrm{km~{}s^{-1}}$. From these two values we adopt $v_{\rm in}$ = -6500 $\mathrm{km~{}s^{-1}}$. From these fits we conclude that the emitting region of the Mg II is well described by a Gaussian profile with a FWHM $\approx$ 7700 $\mathrm{km~{}s^{-1}}$ and an inner radius of $v_{\rm in}$ = -6500 $\mathrm{km~{}s^{-1}}$ that has been attenuated on the red side due to the opacity of the ejecta and cool dense shell. We note that as the profile approaches the peak flux from the blue side, the rise is shallower than the Gaussian profile, possibly indicating asymmetry. Finally, we note that if the CSM is asymmetric, inner and outer shell velocities would represent projected velocities.
@@
Narrow absorption lines are present in the spectra from Fe II $\lambda 2344$, $\lambda 2374$, $\lambda 2383$, $\lambda 2587$, $\lambda 2600$ and Mg II $\lambda\lambda 2796,2802$, $\lambda 2852.96$. These lines are centered at the host redshift and there is no evolution in their equivalent widths with time. From this we conclude that these lines are either from the interstellar medium (ISM) in the host galaxy or distant CSM and remove them from the remainder of the analysis.

### v2

@@ Abstract
SN 2023ixf was discovered in M101 within a day of explosion and rapidly classified as a Type II supernova with flash features. Here we present ultraviolet (UV) spectra obtained with the Hubble Space Telescope 14, 19, 24, and 66 days after explosion. Interaction between the supernova ejecta and circumstellar material (CSM) is seen in the UV throughout our observations in the flux of the first three epochs and asymmetric Mg II emission on day 66. We compare our observations to CMFGEN supernova models that include CSM interaction ($\dot{M}<10^{-3}$ $M\mathrm{{}_{\odot}~{}yr^{-1}}$) and find that the power from CSM interaction is decreasing with time, from $L_{\rm sh}\approx 5\times 10^{42}$ $\mathrm{erg~{}s^{-1}}$ to $L_{\rm sh}\approx 1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$ between days 14 and 66. We examine the contribution of individual atomic species to the spectra on days 14 and 19, showing that the majority of the features are dominated by iron, nickel, magnesium, and chromium absorption in the ejecta. The UV spectral energy distribution of SN 2023ixf sits between that of supernovae which show no definitive signs of CSM interaction and those with persistent signatures assuming the same progenitor radius and metallicity. Finally, we show that the evolution and asymmetric shape of the Mg II $\lambda\lambda$2796, 2802 emission are not unique to SN 2023ixf. These observations add to the early measurements of dense, confined CSM interaction, tracing the mass-loss history of SN 2023ixf to $\sim 33$ yr prior to the explosion and the density profile to a radius of $\sim 5.7\times 10^{15}$ cm. They show the relatively short evolution from a quiescent red supergiant wind to high mass loss.
Type II supernovae; Red supergiant stars; Stellar mass loss; Ultraviolet transient source; Ultraviolet spectroscopy; Hubble Space Telescope
@@ 1 Introduction
Hydrogen-rich supernovae (Type II) are thought to come from red supergiants (RSGs) with zero-age main sequence masses between $\sim 8$ and $\sim 25$ $M_{\mathrm{\odot}}$. While it is accepted that all RSGs lose mass, there are no quantitative mass-loss rates derived from stellar evolution theory. Adding to this, there is no consensus about the empirical mass-loss rates for RSGs; Massey et al. (2023) infer mass-loss rates higher than the canonical rates of de Jager et al. (1988) from the RSG luminosity function, while other studies (e.g. Beasor et al., 2020), measure lower mass-loss rates from the infrared (IR) excess of RSGs. Different mass-loss rates affect the final star that explodes, impacting the final hydrogen envelope mass as well as the density structure of the circumstellar material (CSM) into which the supernova explodes.
@@
Observational evidence supporting these predictions is unclear. Sample analyses found contradictory results about whether supernovae with flash features are systematically bluer and more luminous than noninteracitng supernovae (Khazov et al., 2016; Bruch et al., 2023). Additional contradictions are found in individual objects (e.g. SN 2018zd, Hiramatsu et al. 2021; SN 2020pni, Terreran et al. 2022; various, Jacobson-Galán et al. 2024; PTF11iqb, Smith et al. 2015; SN 2013fs, Bullivant et al. 2018). Spectroscopically, while an absorption feature is often observed blueward of $\mathrm{H\alpha}$ (Gutiérrez et al., 2017) and has been interpreted as high-velocity hydrogen due to CSM interaction, the relationship between the sample of supernovae with this feature and those with flash features is unclear. UV observations of Type II supernovae are especially sparse, with early-time data limited by the availability of sensitive rapid-response UV telescopes, and with late-time observations limited by the fading of the UV when CSM interaction is not present (Panagia et al., 1980; Pettini et al., 1982; Cassatella, 1987; Baron et al., 2000; Fransson et al., 2005; Vasylyev et al., 2022, 2023a).
@@
The study of Type II supernovae with dense, confined CSM was recently enriched by the explosion of SN 2023ixf in the nearby galaxy M101. The proximity, early discovery and classification (Itagaki, 2023; Perley & Gal-Yam, 2023), and identification of narrow emission lines have led to this being the best-observed supernova with flash features to date. Of particular note is the periodic pre-explosion light curve of the RSG progenitor in the optical and IR (Jencson et al., 2023; Kilpatrick et al., 2023; Niu et al., 2023; Qin et al., 2023; Soraisam et al., 2023; Xiang et al., 2024), the incredibly deep multiband pre-discovery data which characterize the shock breakout and light-curve rise starting at an absolute magnitude of $V\approx-11$ together with high-cadence photometry through the end of the plateau phase (Jacobson-Galán et al., 2023; Hiramatsu et al., 2023; Hosseinzadeh et al., 2023; Sgro et al., 2023; Teja et al., 2023; Yamanaka et al., 2023; Zimmerman et al., 2024; Bersten et al., 2024; Li et al., 2024; Martinez et al., 2024), intranight optical spectroscopy for more than a week after discovery (Bostroem et al., 2023a; Jacobson-Galán et al., 2023; Teja et al., 2023; Hiramatsu et al., 2023; Yamanaka et al., 2023; Zhang et al., 2023), early-time optical spectropolarimetry (Vasylyev et al., 2023b; Singh et al., 2024), high-resolution spectroscopy of the first 17 days of evolution (Smith et al., 2023; Hiramatsu et al., 2023; Zimmerman et al., 2024), X-ray and radio detections (Berger et al., 2023; Grefenstette et al., 2023; Matthews et al., 2023; Chandra et al., 2024), and the first early-time UV spectra of a supernova with flash features (Zimmerman et al., 2024). While a coherent picture of these observations is still being explored, they point toward a confined, dense CSM with a radius $R\lesssim 10^{15}$ cm and a mass-loss rate $\dot{M}\approx 10^{-2}$$10^{-4}$ $M_{\mathrm{\odot~{}yr^{-1}}}$.
@@ 2 Observations and Data Reduction
HST observations were obtained on days 14, 19, 24, and 66 with the CCD detector of the Space Telescope Imaging Spectrograph (STIS) using the 52″$\times$ 0.″2 slit at the E1 position to mitigate flux loss to charge transfer efficiency loss and the G230LB, G430L, and G750L gratings (DD GO-17313; Bostroem et al., 2023b). Unfortunately, our planned visit on day 50 failed to acquire guide stars. However, the observation was successfully repeated on day 66. We obtained at least four exposures with each grating to automatically remove cosmic rays. A list of observations is given in Table 1.
@@
The G750L grating suffers fringing at the reddest wavelengths, which can be corrected using a contemporaneous fringe flat. We obtained a fringe flat with each visit using the $0.\arcsec 3\times 0.\arcsec 09$ aperture, as recommended for the E1 aperture position. The fringe flat was applied to the G750L observations using the module stistools.defringe and 1D spectra were extracted using the stistools.x1d routine.
We utilized data-quality flags 16 (high dark rate) and 512 (bad reference pixel) to eliminate bad pixels. In particular, we note that narrow emission features that appear in all spectra are flagged as high dark rate pixels and therefore removed from the analysis. We confirm that these are isolated pixels with high count rates in the dark files by examining the extraction location in the two-dimensional dark images themselves. Nevertheless, it is possible that there is a real unresolved emission line at the same pixel as at least one of these features. However, this would be contaminated by the high dark rate and therefore unable to provide any constraint on the CSM. Thus, given the characteristics described above, we treat these features as artifacts and do not consider them further.
@@
We add to this dataset the HST observations taken between days 3 and 11 as part of GO-17205 (Zimmerman et al., 2022), which have been presented by Zimmerman et al. (2024). 1D-extracted and flux-calibrated files were downloaded from the Mikulski Archive for Space Telescopes. As noted by Zimmerman et al. (2024), some of these exposures suffered from pointing, acquisition, and/or saturation issues, which led to unreliable flux due to the supernova not being aligned in the slit and the nonlinearity of the CCD response near the saturation point. To mitigate these issues, we scale all observations with each grating in a given visit to the spectrum with the most flux (assuming this is the best-centered observation) using a first-degree polynomial fit to the ratio of the spectra as the scale factor. A constant multiple of a few percent was then applied to align the G430L and G750L observations to the G230LB observation, and then the fluxes were combined by taking the median value at each wavelength, excluding pixels with data-quality flags of 16 and 512. We note that the majority of pixels are saturated in the range $\sim 3200$–5000 Å in visit 1, $\sim 2100$–7600 Å in visits 2 and 3, and $\sim 2900$–7900 Å in visit 4, making these wavelength ranges the most unreliable in terms of flux calibration. However, in this paper, we are primarily analyzing the later epochs ($14-66$ days) which are not saturated and only use the early epochs ($<10$ days) to identify the presence and timing of relevant spectral features. Thus, we do not attempt to further correct for this effect.
@@ 3 Spectral Evolution
The near-UV (NUV) through NIR evolution of SN 2023ixf from day 3 to 66 is shown in Figure 1. No Type II supernovae (excluding Type IIn supernovae) since SN 1987A have been observed this early in the UV and with this cadence, revealing unprecedented details of the evolution. As has been noted by other authors, narrow emission lines are present in the early-time spectra, including notably a narrow P Cygni line from N IV $\lambda 1718$ which is present in the first spectrum and fades by day 8, and C III $\lambda 2297$ which is present from day 4 to day 8 (Yamanaka et al., 2023; Jacobson-Galán et al., 2023; Bostroem et al., 2023a; Teja et al., 2023; Hiramatsu et al., 2023; Zimmerman et al., 2024). These early narrow lines indicate dense CSM surrounding the supernova that has not been shocked. By day 14, these lines disappear and the optical spectrum is that of a young Type II supernova. Although slow to develop as a result of early interaction, a typical Type II supernova optical spectrum emerges with prominent Doppler-broadened P Cygni profiles in H I, He I, and other metals.
@@
Figure 1: Evolution of the NUV to NIR spectra of SN 2023ixf from 3.5 to 66 days after explosion. Prominent features (N IV: 1718, 7109, 7122 Å; C III: 2297 Å; Mg II: 2796, 2802 Å; He II: 3203, 4685.5, 4860 Å; C IV: 5801, 5812 Å; $\mathrm{H\alpha}$: 6563 Å; He I: 6678.1 Å; Ca II: 8498, 8542, 8662 Å) are marked with vertical lines at their rest wavelength and labeled at the top of the figure. Narrow ISM absorption lines have been removed to highlight the supernova spectra.
@@
A more detailed view of the evolution from day 14 through day 66 is shown in Figure 2. Although still dominated by iron line blanketing (the broad absorption of flux due to a forest of iron lines that blend together), the UV continuum flux is clearly present in the day 14 spectrum. This is unusual for classical Type II supernovae in which the UV flux fades rapidly (Baron et al., 2000; Brown et al., 2007; Vasylyev et al., 2023a; Bostroem et al., 2023c).
@@
Figure 2: A comparison of the spectra of SN 2023ixf from 14.25 to 66.25 days post-explosion (solid lines) with the best-matching CSM interaction model (semi-transparent). The zero flux levels for each epoch are drawn as a dot-dashed line in the same color as the observed spectrum at that epoch. Over time, the UV emission declines and P Cygni features develop in the optical, until day 66, when only Mg II $\lambda\lambda$2796, 2802 is visible in the UV. This corresponds to a sequence of models with decreasing interaction power. Even though the model spectra do not match the SED of the observations, especially in the first two spectra where they appear to be cooler (resulting in lower-ionization species and more UV line blanketing), the match is impressive given the 1D nature and the ad-hoc implementation of the mass, density, and clumping of the cool dense shell. Prominent spectral features have been marked with dotted lines at 7000 $\mathrm{km~{}s^{-1}}$, the approximate location of the absorption throughout the spectroscopic sequence. The “emission” feature at 1925 Å is marked at rest with a dashed line. Narrow ISM absorption lines have been removed to better highlight the supernova spectra. Observed and model spectra at each epoch have been placed at the same distance with the same offset applied.
@@
A comparison of the Mg II and $\mathrm{H\alpha}$ profiles provides insight into the origin of the Mg II emission (see Figure 3). When CSM is present, a cool dense shell forms at the interface between the ejecta and CSM. While the $\mathrm{H\alpha}$ emission originates from the ejecta and the cool dense shell, the Mg II is thought to originate only from the shell and should thus be at a similar velocity as the outer edges of the ejecta. We compare Mg II and $\mathrm{H\alpha}$ in Figure 3, finding that the profiles extend to similar velocities on both the red and blue sides, confirming that the Mg II is emitted from the cool dense shell at the outer edge of the ejecta. However, $\mathrm{H\alpha}$ sits in the optical where there is continuum and forms throughout the ejecta and cool dense shell, leading to a P Cygni profile, while Mg II, without UV continuum and forming in the outer ejecta, is seen solely in emission. There is a shallow, narrow dip at the blue edge of the $\mathrm{H\alpha}$ profile that could be absorption from the cool dense shell at $\sim 9200$ $\mathrm{km~{}s^{-1}}$, although it is not clear if this feature is real given the signal-to-noise ratio (S/N) of the spectrum.
@@
Figure 3: Left: $\mathrm{H\alpha}$ and Mg II line profiles on day 66 in velocity space of SN 2023ixf compared to the CSM interaction models of Dessart & Hillier (2022) with $L_{\rm sh}=1\times 10^{40}$ $\mathrm{erg~{}s^{-1}}$, scaled to the distance of SN 2023ixf. The shapes of both lines are well captured by the models, given that the cool dense shell is at a higher velocity in the models than in SN 2023ixf. Right: Mg II profile (black) fit with a pure Gaussian (pink) modeling spherical ejecta, a symmetric shell (blue dotted) modeling emission from the cool dense shell in the absence of optical-depth effects, and an asymmetric shell (green dashed) modeling the cool dense shell with optical depth.
@@
To characterize the Mg II feature further, we fit three different profiles in Figure 3, representing different physical configurations. A spherical distribution of material with a Gaussian emitting profile will produce a Gaussian profile. Although the profile is clearly not Gaussian, it could be coming from a spherical distribution of material that is being attenuated. To explore the nonattenuated spherical distribution, we fit a Gaussian to the blue side of the Mg II profile from $-10,000$ to $-7000$ $\mathrm{km~{}s^{-1}}$. This fit is shown as a pink line in Figure 3; it has a mean of $\mu=0$ $\mathrm{km~{}s^{-1}}$ and a FWHM of 7780 $\mathrm{km~{}s^{-1}}$. The observed blue side of Mg II is well described by a Gaussian profile. An alternate emitting region could be a sphere with a Gaussian profile that has had the center removed to form a thick shell. This configuration will have a flat-topped profile with Gaussian wings (Jerkstrand, 2017) if the cool dense shell and ejecta are optically thin. Given our slanted top profile, we use the maximum of our observed flux as the flat-topped maximum of the model profile. This defines the inner radius velocity of the shell as $v_{\rm in}=-6800$ $\mathrm{km~{}s^{-1}}$, and the profile is shown as a dotted-blue line in the right panel of Figure 3.
@@
To model the slated top of the profile, we model the line as if it were generated by an asymmetric shell of material (i.e., a shell in which there is more flux from one side than the other reaching the observer). Following Kwok et al. (2023, 2024), we model this asymmetric shell as a Gaussian profile with an off-center hole cut out, which results in a slanted top instead of a flat-topped profile. We parameterize this model by the FWHM and mean of the Gaussian, the velocity of the center of the hole ($v_{\rm c}$), and the velocity of the radius of the hole ($v_{\rm in}$). We find FWHM = 7580 $\mathrm{km~{}s^{-1}}$, $\mu=-2995$ $\mathrm{km~{}s^{-1}}$, $v_{\rm in}=5000$ $\mathrm{km~{}s^{-1}}$, and $v_{\rm c}=1750$ $\mathrm{km~{}s^{-1}}$, where $v_{\rm c}$ and $v_{\rm in}$ are measured relative to $\mu$. This model fits the data exceptionally well and is shown as a green dashed line in the right panel of Figure 3. While the CSM is likely asymmetric, we caution that this parameterization is more likely to indicate different optical depths along the line of sight through the ejecta and cool dense shell, and should not be taken as a literal description of the ejecta. It does, however, demonstrate that this profile can be created by a shell of material in which a higher flux of photons from the blue side of the ejecta is reaching us, than photons from the red side. Additionally, it defines the inner edge of the blue side of the shell to be at velocity $-6245$ $\mathrm{km~{}s^{-1}}$(where the profile peaks), close to our earlier estimate of $-6800$ $\mathrm{km~{}s^{-1}}$. From these two values, we adopt $v_{\rm in}$ = -6500 $\mathrm{km~{}s^{-1}}$. From these fits we conclude that the emitting region of the Mg II is well described by a Gaussian profile with a FWHM $\approx$ 7700 $\mathrm{km~{}s^{-1}}$ and an inner radius of $v_{\rm in}$ = -6500 $\mathrm{km~{}s^{-1}}$ that has been attenuated on the red side due to the opacity of the ejecta and cool dense shell. We note that as the profile approaches the peak flux from the blue side, the rise is shallower than the Gaussian profile, possibly indicating asymmetry. Finally, we note that if the CSM is asymmetric, inner and outer shell velocities would represent projected velocities.
@@
Narrow absorption lines are present in the spectra from Fe II $\lambda 2344$, $\lambda 2374$, $\lambda 2383$, $\lambda 2587$, $\lambda 2600$ and Mg II $\lambda\lambda 2796,2802$, $\lambda 2852.96$. These lines are centered at the host redshift and there is no evolution in their equivalent widths with time. From this, we conclude that these lines are either from the interstellar medium (ISM) in the host galaxy or distant CSM and remove them from the remainder of the analysis.